This question seems to have me stummped, Im not sure exactly sure how to correctly solve it, is there any out there that can help me tackle this task?
The question is 8x-4y=16
y=2x-4
Thank your for your kindness and generosity in advance.

Answer by ankor@dixie-net.com(22740)

This equation has no unique solution as we shall see when we try to solve it, however, here is the method that may help you
The question is:
8x -4y = 16
:
y = 2x-4; see this equation here, it means that anywhere you have a "y" you can substitute (2x-4) for it..
:
With this in mind, take the 1st equation: 8x - 4y = 16 and substitute (2x-4) for y. Then you have have:
8x - 4(2x-4) = 16; notice we only have 1 unknown (x) now.
:
Multiply what's inside the brackets; remember,a negative outside the brackets means you must change the signs inside the brackets.
8x - 8x + 16 = 16
:
Notice that 8x - 8x = 0 so we end up with 16 = 16 which of course is true but any value of x will be true also. I think this is called an identity, no unique solution. That leaves the system unsolved.
:
Hope this helps you